OpenStudy (anonymous):

Which ratios form a proportion? 4/9,12/25 3/7, 18/42 5/11, 20/45 5/8, 15/25 PLEASE HELP. MY GRADE DEPENDS ON IT.

jimthompson5910 (jim_thompson5910):

4/9 = 12/25 4*25 = 9*12 ... cross multiply 100 = 108 ... which is false so 4/9 = 12/25 is not true and 4/9,12/25 don't form a proportion

jimthompson5910 (jim_thompson5910):

do the same with the rest